1. Three Identical lamps are connected in series to a 12 volt battery. What is the voltage drop across each lamp.?
  2. Three 50- Ω resistors are connected in series . The equivalent resistance of the combination is?
  3. Two resistors, R1 = 2.0 kΩ and R2= 3.0kΩ are connected in parallel and their combination is connected in series to a fully charged, 150 µF capacitor. When the switch is opened the capacitor begins to discharge. What is the time constant for the discharge?
  4. Four 50 Ω resistor are connected in parallel. The equivalent resistance of the combination is?
  5. How many 4Ω capacitor must be connected in parallel to have a total resistance of 0.8 Ω ?
1
Expert's answer
2021-06-09T08:05:56-0400

1. The voltage drop across each lamp is


"V_L=\frac V3=4\text V."

2. The equivalent resistance of the combination is


"R_e=R_1+R_2+R_3=50+50+50=150\Omega."

3. The equivalent resistance:


"\frac{1}{R_e}=\frac{1}{R_1}+\frac{1}{R_2},\\\space\\
R_e=\frac{R_1R_2}{R_1+R_2}."

Time constant:


"\tau=RC=\frac{R_1R_2}{R_1+R_2}C=4500000\text s."

4. The equivalent is

"\frac{1}{R_e}=\frac{1}{R_1}+\frac{1}{R_2}+\frac{1}{R_3}+\frac{1}{R_4},\\\space\\
\frac{1}{R_e}=\frac{1}{R}+\frac{1}{R}+\frac{1}{R}+\frac{1}{R},\\\space\\
R_e=\frac R4=12.5\Omega."

5. From the last equation above it follows that


"n=\frac{R}{R_e}=\frac{4}{0.8}=5."